ABC Company sold the rights to use one of their patented processes that will result in them receiving cash payments of $20,000 at the end of every six months for each of the next five years and a lump sum payment of $4,000 at the end of the sixth year. Calculate the total present value of these payments if the interest rate is 12% compounded semi-annually. Use the time value of money factors to answer this question
Answer #1

Given in question:

Time - 5 years

Compounding semi annually, so

N = 5*2 = 10

Discount rate will be = 12/2 = 6%

So 5 years compounding factor of the rate of 6 % for (1/1.06)*10 =9.434

So 6th year NPV factor = (1/1.12)6 =0 .705 ( Here 6 refers power of whole bracket 6)

So finally calculation of NPV :

200000*9.434 = 188680

4000*.705 = 2820

So Total NPV = 191500   ( Answer)